a chemist must dilute 45.8 ml of 1.26 m aqueous barium acetate (Ba(C2H3O2)2) solution until the concentration falls to 1.00 M. he’ll do this by adding distilled water to the solution until it reaches a certain final volume. calculate this final volume, in milliliters. round your answer to 3 significant digits.
1
Expert's answer
2020-02-21T07:50:38-0500

The number of moles of barium acetate in the initial solution equals:

n = c1 × V1,

where n - number of moles, c1 - molar concentration of the initial solution, V1 - volume of the initial solution.

As number of moles is not changed after the dilution, volume of the final solution is:

V2 = n / c2,

where V2 - volume of the final solution, c2 - molar concentration of the final solution, n - number of moles.

As n = c1 × V1, than:

V2 = c1 × V1 / c2 = 1.26 M × 45.8 ml / 1.00 M = 57.7 ml


Answer: 57.7 ml